# PASER [Uncaptioned image]: Post-Training Data Selection for Efficient Pruned Large Language Model Recovery

### v1

@@ Appendix F Extended Experiments on Recovery Training with Knowledge Distillation
Appendix G Detailed Ablation Study Results
In this section, we present comprehensive ablation results of the three key components in PASER: semantic-structural recovery instruction clustering (S2RIC), capability degradation-aware instruction selection (CDAIS), and negative transfer mitigation (NTM). Table 11 shows the detailed performance across different evaluation metrics.
The detailed results reveal the distinct contributions of each component under different pruning schemes. For structured pruning like LLM-Pruner, removing S2RIC leads to significant degradation in both language modeling (perplexity increases from 16.40 to 18.73 on WikiText2) and reasoning tasks (average score drops by 1.43 points), highlighting its importance in addressing uneven capability degradation. The impact of CDAIS is particularly evident under SliceGPT, where its removal causes a 1.63-point drop in average reasoning performance while maintaining relatively stable language modeling metrics, demonstrating its effectiveness in balancing recovery priorities. Under semi-structured pruning (Wanda), all three components show more balanced contributions, with performance drops ranging from 0.68 to 0.83 points when each is removed. This suggests that semi-structured pruning requires a more holistic recovery approach. For unstructured pruning (SparseGPT) where capability degradation tends to be more uniform, NTM plays a particularly crucial role - its removal leads to the largest drop in language modeling performance (perplexity increases from 13.33 to 15.91 on WikiText2) and affects complex reasoning tasks like WinoGrande and ARC-e significantly. Notably, the full PASER framework consistently achieves the best performance across almost all metrics under various pruning schemes, with only occasional exceptions in individual tasks (e.g., OBQA in LLM-Pruner and PIQA in SparseGPT). This comprehensive superiority validates our design choice of combining these three components for effective pruned model recovery.
@@ Appendix G Detailed Ablation Study Results
Appendix H Exploration on Other Possible Clustering Methods
To discuss the impact of different instruction tuning data clustering approaches, we replace our Semantic-structural Recovery Instruction Clustering (S2RIC) module with some other common text clustering method: NMF_TFIDF, LDA_TFIDF, KMeans_TFIDF, Spectral_MTEB, Spectral_BERT (Xu et al., 2024a). The reasoning performance comparison among different PASER versions with such clustering methods is provided in Table 12. From the table, we can find that integrating other instruction clustering methods with PASER can bring the performance decline to some extent among all four pruning schemes. Especially, the clustering method with traditional statistics-based text representation technique, TFIDF, generally behaves worse than semantic embedding-based text representation techniques like BERT. Therefore, we can conclude that our semantic-structural recovery instruction clustering is at least a competitive approach as the clustering component of PASER. Though, comparing these results with those in Table 1, we can observe the advantages of PASER over other general instruction tuning data selection methods can still be stably maintained. This further demonstrates that the potential of the clustering-based data selection for effective and balanced LLM capability recovery.
@@ Appendix H Exploration on Other Possible Clustering Methods
PASER 13.33 23.77 74.79 78.38 66.62 69.03 72.57 38.70 39.40 62.78
Appendix I Case Study for Recovery Instruction Clustering
@@
I.1 Cluster 1: Basic Factual Knowledge and Information Retrieval
@@
I.2 Cluster 2: Language Understanding and Translation
@@
  • Instruction: “Difference in meaning between "done freely" and "freely done¨? For instance, is there any difference in meaning between these two sentences?”’
@@
I.3 Cluster 3: Logical Reasoning and Problem Solving
@@
I.4 Cluster 4: Creative Writing and Text Generation
@@
I.5 Cluster 5: Summarization and Information Extraction
@@
  • Instruction: “Provide a brief summary of the article "A Brief History of the Automobile Industry.”
@@
Appendix J Case Study for Negative Transfer Mitigation
To illustrate the negative transfer mitigation process in PASER, we present the following case study, including both initial concept consistency graph construction and new sample evaluation.
J.1 Constructing Initial Concept Consistency Graph
@@
J.2 Evaluating New Samples
We now evaluate two new samples using PASER’s negative transfer mitigation process.
J.2.1 Sample 1: Accepted
@@
Applying PASER’s negative transfer mitigation process:
@@
Applying PASER’s negative transfer mitigation process:
@@
J.3 Conclusion
This case study demonstrates PASER’s negative transfer mitigation process in action. By accepting Sample 1, PASER allows for the introduction of new, relevant concepts that expand the concept consistency graph without introducing conflicts. By rejecting Sample 2, PASER prevents the introduction of potentially misleading relationships between existing concepts from different domains, thus mitigating the risk of negative transfer during the recovery process.
Appendix K Implementation Details
Most of experiments are conducted on the server with 8 $\times$ RTX 6000 Ada GPUs. During the recovery post-training phase, we take the the low-rank approximation, LoRA (Hu et al., ), to improve the efficiency. The corresponding hyperparameters are set as following: rank=8, batch size=64, epochs=2, learning rate = 1e-4 (Alpaca series experiments), 5e-5 (LaMini series experiments). As for the structured pruning, we set the pruning ratio as 25% for LLaMA2-7B/LLaMA3-8B/Baichuan2-7B and 50% for LLaMA2-13B/LLaMA2-70B/Baichuan-13B models. For the other two kinds of pruning schemes, we follow the previous work (Frantar & Alistarh, 2023); Specifically, we adopt the 2:4 semi-structured sparsity patterns and implement 50% unstructured weight sparsity. Except the experiments for recovery post-training efficiency analysis, we set the ratio of recovery data budget $B$ to original dataset size $N$ as 20% for Alpaca and 4% for LaMini. As for the implementation of concept extraction in Section 3.4, we use the open-source library rake-nltk  [^2] https://pypi.org/project/rake-nltk/. To ensure statistical robustness, all the results reported in this paper are the averages of five runs with different seeds. Statistical significance is also assessed using two-tailed independent t-tests, with results considered significant when $p<0.01$. For facilitating the reproduction of our work, we provide the code in the supplementary materials, also seen in anonymous github https://github.com/BokwaiHo/PASER.
Appendix L Limitation Analysis
@@
Limitations of Concept Consistency Graph. In addition to the above analysis, we further discuss the concept consistency graph’s potential limitations. Indeed, while CCG helps mitigate negative transfer, we acknowledge there are scenarios where semantic conflicts might not be fully captured:
@@
Appendix M Ethics Statement

### v2

@@ Appendix G Extended Experiments on Recovery Training with Knowledge Distillation
Table 11: Recovery performance of different instruction tuning data selection methods on several more recent and advanced LLMs under LLM-Pruner scheme. The ‘bold’ represents the best performance on the same target LLM. Here, the Alpaca is taken as the original dataset. The “Reason” indicates the averaged performance on 7 common sense reasoning datasets.
LLaMA3.1-8B ratio=25% WikiText2$\downarrow$ 6.63 14.85 9.74 12.16 13.84 11.32 10.81 7.56
[…]
Qwen2.5-7B ratio=25% WikiText2$\downarrow$ 6.94 15.82 11.68 13.94 15.37 10.95 10.23 8.16
[…]
Qwen3-8B ratio=25% WikiText2$\downarrow$ 6.68 14.29 9.18 11.83 13.26 10.17 9.45 7.24
[…]
Mixtral-8x7B ratio=25% WikiText2$\downarrow$ 7.52 15.87 12.83 14.32 14.27 13.65 12.98 9.62
[…]
Appendix H Experiments on More Recent LLMs
In addition to the previous experiments conducted on LLaMA2/3 and Baichuan2 families, we also supplement some results on more recent and advanced LLMs: LLaMA3.1-8B  [^2] https://huggingface.co/meta-llama/Llama-3.1-8B, Qwen2.5-7B [Yang et al., 2024]  [^3] https://huggingface.co/Qwen/Qwen2.5-7B, Qwen3-8B  [^4] Qwen/Qwen3-8B. Besides, the Mixture of Experts(MOE) model Mixtral-8$\times$7B [Jiang et al., 2024]  [^5] https://huggingface.co/mistralai/Mixtral-8x7B-v0.1 is also employed to conduct the experiments. The evaluation results have been provided in Table 11. From Table 11, we can observe that PASER consistently outperforms all baseline methods across different recent and advanced LLMs, including the MOE architecture (Mixtral-8x7B). For LLaMA3.1-8B, PASER achieves perplexity scores (7.56 on WikiText2, 17.83 on PTB) that are remarkably close to those of the unpruned model (6.63 and 16.72, respectively), with only a 0.76 point gap in reasoning performance (70.05 vs. 70.81). Notably, PASER demonstrates robust performance across different model architectures. For the MoE-based Mixtral-8x7B, PASER significantly narrows the performance gap from 10.48 points (w/o Training: 61.56 vs. w/o pruning: 72.04) to just 1.17 points (PASER: 70.87 vs. w/o pruning: 72.04), representing a 88.8% recovery of lost capability. This demonstrates that our capability-aware selection approach effectively addresses the specialized recovery needs of sparse models with expert-based architectures. Among the most recent models, Qwen3-8B shows the strongest recovery with PASER, achieving 98.4% of the unpruned model’s reasoning performance (70.58 vs. 71.73). Meanwhile, Qwen2.5-7B demonstrates slightly lower recovery effectiveness, potentially due to differences in its architecture and pretraining approach. These results further validate PASER’s applicability across diverse and state-of-the-art LLM architectures, confirming that our capability degradation-aware instruction selection and negative tuning effects mitigation strategies generalize well beyond the models presented in the main paper.
@@ Appendix H Experiments on More Recent LLMs
Appendix I Detailed Ablation Study Results
In this section, we present comprehensive ablation results of the three key components in PASER: semantic-structural recovery instruction clustering (S2RIC), capability degradation-aware instruction selection (CDAIS), and negative tuning effects mitigation (NTEM). Table 12 shows the detailed performance across different evaluation metrics.
The detailed results reveal the distinct contributions of each component under different pruning schemes. For structured pruning like LLM-Pruner, removing S2RIC leads to significant degradation in both language modeling (perplexity increases from 16.40 to 18.73 on WikiText2) and reasoning tasks (average score drops by 1.43 points), highlighting its importance in addressing uneven capability degradation. The impact of CDAIS is particularly evident under SliceGPT, where its removal causes a 1.63-point drop in average reasoning performance while maintaining relatively stable language modeling metrics, demonstrating its effectiveness in balancing recovery priorities. Under semi-structured pruning (Wanda), all three components show more balanced contributions, with performance drops ranging from 0.68 to 0.83 points when each is removed. This suggests that semi-structured pruning requires a more holistic recovery approach. For unstructured pruning (SparseGPT) where capability degradation tends to be more uniform, NTEM plays a particularly crucial role - its removal leads to the largest drop in language modeling performance (perplexity increases from 13.33 to 15.91 on WikiText2) and affects complex reasoning tasks like WinoGrande and ARC-e significantly. Notably, the full PASER framework consistently achieves the best performance across almost all metrics under various pruning schemes, with only occasional exceptions in individual tasks (e.g., OBQA in LLM-Pruner and PIQA in SparseGPT). This comprehensive superiority validates our design choice of combining these three components for effective pruned model recovery.
Appendix J Hyperparameter Robustness Analysis
To demonstrate the robustness of PASER to various hyperparameter settings, we conducted a comprehensive sensitivity analysis. This analysis helps establish that our method’s strong performance is not dependent on specific hyperparameter configurations but remains consistent across a reasonable range of settings.
J.1 Sensitivity to Embedding Dimension
The dimensionality of the manifold representation after applying the diffusion kernel is controlled by parameter $d$. We evaluate PASER’s performance with $d$ ranging from 4 to 64 on LLaMA2-7B under LLM-Pruner (ratio=25%). As shown in left subfigure of Figure 5, clustering results remains consistent (Rand Index $>$0.85) across this range, and average reasoning performance variations is controlled within $\pm$0.2 points. This indicates that PASER is relatively stable to the specific choice of embedding dimension, though we select $d=16$ based on the eigenvalue decay pattern, where eigenvalues beyond this dimension contribute negligibly to the representation.
Figure 5: (a) Sensitivity to embedding dimension $d$ after manifold learning, here the clustering results under $d=16$ is taken as the reference; (b) Robustness to temperature parameter $\tau$ in Equation 5.
J.2 Robustness to Temperature Parameter
For the Jensen-Shannon divergence (JSD) calculation in the capability degradation score, we investigate the effect of temperature $\tau$ in the softmax function (Equation 5). In right subfigure of Figure 5, we test values in the range $\tau\in\{0.5,0.75,1.0,1.25,1.5\}$ and observe that while lower temperatures ($\tau<0.75$) lead to slightly sharper capability distinctions, the overall performance differences are minimal ($<$0.3 points). Our default setting of $\tau=1.0$ provides a good balance between sensitivity to capability differences and stability across pruning schemes.
Overall, our hyperparameter sensitivity analysis demonstrates that PASER’s performance is robust across a wide range of hyperparameter settings. This stability enhances the practical applicability of our method, as it does not require extensive hyperparameter tuning to achieve strong performance in real-world scenarios.
Appendix K Empirical Exploration for Component Selection
To demonstrate the selection optimality of each component in our PASER framework, we conduct the empirical exploration on alternative initial instruction embedding models, dimensionality reduction methods, clustering methods, divergence measurement methods, and budget allocation strategy, to investigate if they can achieve better final recovery performance.
Table 13: Recovery performance of multiple PASER versions integrated with initial instruction embedding models under various pruning schemes on LLaMA2-7B model. The PASER(SentenceBERT) is the version we employ in the main body. The ‘bold’ represents the best performance under the same pruning scheme. Here, the Alpaca is taken as the original dataset.
Pruning
Recovery
Post-training
[…]
PASER(Qwen2.5-7B) 16.38 26.33 67.27 77.32 69.01 67.00 67.86 39.56 39.80 61.12
PASER(LLaMA3-8B) 16.37 26.31 67.29 77.32 69.01 67.00 67.84 39.58 40.00 61.15
PASER(SentenceBERT) 16.40 26.35 67.25 77.29 68.98 66.97 67.84 39.54 39.80 61.10
PASER(Qwen2.5-7B) 12.22 21.51 72.78 79.86 73.94 69.20 71.39 41.86 41.30 64.33
PASER(LLaMA3-8B) 12.21 21.50 72.80 79.86 73.94 69.20 71.40 41.85 41.30 64.34
PASER(SentenceBERT) 12.24 21.53 72.75 79.84 73.92 69.18 71.37 41.82 41.30 64.31
PASER(Qwen2.5-7B) 14.11 27.20 70.79 77.87 71.78 66.26 68.32 39.03 40.10 62.02
PASER(LLaMA3-8B) 14.10 27.22 70.80 77.90 71.80 66.29 68.33 39.07 40.10 62.04
PASER(SentenceBERT) 14.13 27.22 70.77 77.87 71.78 66.26 68.30 39.04 40.10 62.02
PASER(Qwen2.5-7B) 13.62 24.36 74.65 78.26 66.46 68.80 72.50 38.72 39.30 62.67
PASER(LLaMA3-8B) 13.50 23.98 74.72 78.31 66.55 68.96 72.50 38.63 39.30 62.71
PASER(SentenceBERT) 13.33 23.77 74.79 78.38 66.62 69.03 72.57 38.70 39.40 62.78
K.1 Exploration on Other Possible Instruction Embedding Models
In the Semantic-Structal Recovery Instruction Clustering module of our PASER framework, we adopt the SentenceBERT [Reimers and Gurevych, 2019] to obtain the initial instruction embedding, considering that SentenceBERT has demonstrated strong transfer capability across various text semantic tasks. To further validate that SentenceBERT is competitive enough here with the empirical evidence, we replace it with two other advanced LLMs: Qwen2.5-7B [Yang et al., 2024] and LLaMA3-8B [Dubey et al., 2024] here and conduct the experiments. As shown in Table 13, using much larger pretrained language models like Qwen2.5-7B and LLaMA3-8B as embedding generators provides marginal performance improvements (less than 0.05 points on average) compared to SentenceBERT. While LLaMA3-8B achieves the highest performance in three configurations, these gains are negligible given the substantially increased computational costs associated with using such large models for initial embeddings. Besides, under the SparseGPT, sparsity=50% setting, SentenceBERT performs obviously better than LLaMA3-8B and Qwen2.5-7B. This finding validates our design choice of SentenceBERT as the default embedding model for PASER, as it strikes an optimal balance between semantic representation quality and computational efficiency. SentenceBERT requires significantly fewer resources while still enabling effective capability-specific clustering that supports our targeted recovery approach. The results demonstrate that while more powerful embedding models may theoretically provide better semantic representations, in practice, the performance difference is minimal for our specific task of recovery data selection.
Table 14: Recovery performance of multiple PASER versions integrated with different dimensionality reduction approaches under various pruning schemes on LLaMA2-7B model. The PASER(Diffusion) is the version we employ in the main body. The ‘bold’ represents the best performance under the same pruning scheme. Here, the Alpaca is taken as the original dataset.
Pruning
Recovery
Post-training
[…]
PASER(UMAP) 16.92 27.83 66.31 76.84 67.59 65.85 66.92 38.96 39.20 60.31
PASER(PCA) 17.05 28.16 66.18 76.73 67.46 65.73 66.84 38.87 39.10 60.18
PASER(t-SNE) 17.21 28.42 66.05 76.68 67.34 65.64 66.74 38.73 39.00 60.05
PASER(Diffusion) 16.40 26.35 67.25 77.29 68.98 66.97 67.84 39.54 39.80 61.10
PASER(UMAP) 14.83 25.42 71.15 78.91 72.25 67.84 69.95 40.82 40.30 63.03
PASER(PCA) 13.87 23.75 71.73 79.31 72.79 68.48 70.53 41.28 40.70 63.55
PASER(t-SNE) 13.59 23.24 72.04 79.53 73.06 68.73 70.81 41.43 40.90 63.79
PASER(Diffusion) 12.24 21.53 72.75 79.84 73.92 69.18 71.37 41.82 41.30 64.31
PASER(UMAP) 15.84 30.25 69.26 77.42 70.31 65.82 67.84 38.67 39.00 61.19
PASER(PCA) 15.46 29.48 69.14 77.35 70.27 65.74 67.79 38.75 39.60 61.23
PASER(t-SNE) 14.72 28.21 70.14 77.62 71.08 66.01 68.02 38.85 39.80 61.65
PASER(Diffusion) 14.13 27.22 70.77 77.87 71.78 66.26 68.30 39.04 40.10 62.02
PASER(UMAP) 14.89 26.31 73.25 77.45 65.15 68.47 71.28 37.82 38.80 61.75
PASER(PCA) 14.62 25.84 73.51 77.68 65.52 68.73 71.61 38.14 38.90 62.01
PASER(t-SNE) 13.98 24.95 74.13 78.04 66.14 68.86 72.09 38.43 39.10 62.40
PASER(Diffusion) 13.33 23.77 74.79 78.38 66.62 69.03 72.57 38.70 39.40 62.78
K.2 Exploration on Other Possible Dimensionality Reduction Methods
To validate the suitability of diffusion kernel as dimensionality reduction approach which focus on uncovering the instruction geometric structure in the semantic space to facilitate the following capability clustering process, we replace it with other alternative schemes and compare the performance. As shown in Table 14, among different dimensionality reduction methods, our diffusion kernel approach consistently outperforms alternatives across all pruning schemes. The diffusion kernel’s ability to preserve manifold structure while adapting to the intrinsic geometry of instruction data proves particularly advantageous for capability-aware clustering. UMAP, PCA, and t-SNE each demonstrate reasonable performance, but they fall short by 0.38-1.28 points compared to the diffusion kernel approach on average reasoning tasks. The performance gap is especially pronounced under structured pruning schemes like LLM-Pruner (0.79 points advantage) and SliceGPT (0.52 points advantage), where capability degradation tends to be more severe and uneven. This confirms that the diffusion kernel’s non-linear dimensionality reduction better captures the complex relationships between instructions targeting similar capabilities, enabling more effective recovery prioritization. While t-SNE shows competitive performance in some scenarios, its sensitivity to hyperparameters and initialization makes it less stable for diverse instruction sets compared to our approach.
@@ Appendix I Detailed Ablation Study Results
PASER(S2RIC) 13.33 23.77 74.79 78.38 66.62 69.03 72.57 38.70 39.40 62.78
K.3 Exploration on Other Possible Clustering Methods
To discuss the impact of different instruction tuning data clustering approaches, we replace our Semantic-structural Recovery Instruction Clustering (S2RIC) module with some other common text clustering method: NMF_TFIDF, LDA_TFIDF, KMeans_TFIDF, Spectral_MTEB, Spectral_BERT [Xu et al., 2024a]. The reasoning performance comparison among different PASER versions with such clustering methods is provided in Table 15. From the table, we can find that integrating other instruction clustering methods with PASER can bring the performance decline to some extent among all four pruning schemes. Especially, the clustering method with traditional statistics-based text representation technique, TFIDF, generally behaves worse than semantic embedding-based text representation techniques like BERT. Therefore, we can conclude that our semantic-structural recovery instruction clustering is at least a competitive approach as the clustering component of PASER. Though, comparing these results with those in Table 1, we can observe the advantages of PASER over other general instruction tuning data selection methods can still be stably maintained. This further demonstrates that the potential of the clustering-based data selection for effective and balanced LLM capability recovery.
Table 16: Recovery performance of multiple PASER versions integrated with different capability degradation assessment approaches under various pruning schemes on LLaMA2-7B model. The PASER(JSD) is the version we employ in the main body. The ‘bold’ represents the best performance under the same pruning scheme. Here, the Alpaca is taken as the original dataset.
Pruning
Recovery
Post-training
[…]
PASER(KLD) 16.91 27.54 66.52 76.95 68.24 65.84 66.58 38.74 39.70 60.37
PASER(WD) 16.73 27.26 66.81 77.10 68.58 66.18 67.08 39.09 39.80 60.59
PASER(JSD) 16.40 26.35 67.25 77.29 68.98 66.97 67.84 39.54 39.80 61.10
PASER(KLD) 13.81 23.64 71.59 78.87 72.28 67.85 69.94 40.84 40.50 63.12
PASER(WD) 13.10 22.75 72.08 79.25 72.93 68.39 70.48 41.26 40.90 63.61
PASER(JSD) 12.24 21.53 72.75 79.84 73.92 69.18 71.37 41.82 41.30 64.31
PASER(KLD) 15.29 29.08 69.85 77.13 70.65 65.47 67.58 38.52 39.20 61.20
PASER(WD) 14.64 28.21 70.42 77.54 71.23 65.86 67.95 38.78 39.70 61.64
PASER(JSD) 14.13 27.22 70.77 77.87 71.78 66.26 68.30 39.04 40.10 62.02
PASER(KLD) 14.76 25.97 73.51 77.71 65.87 68.35 71.82 38.36 38.90 62.07
PASER(WD) 14.21 25.18 74.15 78.05 66.24 68.68 72.19 38.53 39.10 62.42
PASER(JSD) 13.33 23.77 74.79 78.38 66.62 69.03 72.57 38.70 39.40 62.78
K.4 Exploration on Other Possible Capability Degradation Assessment Methods
To further demonstrate the advantages of utilizing Jensen-Shannon divergence (JSD) as the capability degradation assessment approach (elaborated in Section 3.3), we compare it with alternative schemes like Kullback–Leibler divergence (KLD) [Kullback and Leibler, 1951] and Wasserstein distance (WD) [Vaserstein, 1969]. As shown in Table 16, JSD consistently outperforms alternative divergence measures across all pruning schemes. While KLD is a widely used dissimilarity measure, its asymmetric nature $(\text{KLD}(P||Q)\neq\text{KLD}(Q||P))$ makes it less suitable for comparing output distributions between original and pruned models, resulting in a 0.73 points average performance drop compared to JSD. The WD performs better than KLD, but still falls 0.51 points short of JSD on average reasoning performance. Although WD can capture gradual distribution shifts that might be missed by other divergence measures, it appears to be more sensitive to outlier probability differences that don’t necessarily correlate with capability degradation. This suggests that major shifts in probability mass, rather than fine-grained transportation costs, are more indicative of capability deterioration. JSD’s symmetry and boundedness (always between 0 and 1) provide a more stable and interpretable measure of capability divergence, enabling more accurate prioritization of severely affected capabilities. Its ability to balance sensitivity to both major and minor distribution shifts makes it particularly effective for detecting subtle capability degradations that might be missed by other measures. The superior performance of JSD across all four pruning schemes validates our design choice in Section 3.3.
Table 17: Recovery performance of different budget allocation strategies under various pruning schemes on LLaMA2-7B model. The Linear proportional is the version we employ in the main body. The ‘bold’ represents the best performance under the same pruning scheme. Here, the Alpaca is taken as the original dataset.
Allocation Strategy LLM-Pruner SliceGPT Wanda SparseGPT
Equal 60.34 63.52 61.28 62.05
Square-root scaling 60.68 63.81 61.56 62.31
Logarithmic scaling 60.51 63.65 61.43 62.18
Linear proportional 61.10 64.31 62.02 62.78
K.5 Exploration on Other Possible Budget Allocation Strategies
Beyond our proportional allocation strategy in Equation 6, we explore alternative allocation strategies including equal allocation, square-root scaling ($n_{k}\propto\sqrt{\text{CDS}(c_{k})}$), and logarithmic scaling ($n_{k}\propto\log(1+\text{CDS}(c_{k}))$). As shown in Table 17, the linear proportional allocation consistently outperforms alternatives, showing $\sim$0.4-0.8 points higher average performance across pruning schemes. This validates our design choice while demonstrating that PASER maintains reasonable performance even with different allocation strategies.
Appendix L Fine-grained Data Selection Time Consumption Analysis
In addition to the recovery post-training efficiency analysis provided in the Figure 2 and Section 4.2, we also analyze the fine-grained time consumption of each component to provide a deeper understanding for PASER’s efficiency advantages during data selection phase.
Table 18: Time consumption breakdown of PASER components on different-sized datasets.
Component Alpaca (52K) LaMini (2.58M)
Time (min) Percentage Time (hrs) Percentage
Semantic-Structural Recovery Instruction Clustering
SentenceBERT Embedding 0.63 11% 0.12 10%
Diffusion Kernel 1.64 28% 0.36 31%
NMF Spectral Clustering 0.82 14% 0.15 13%
Capability Degradation-Aware Instruction Selection
JSD Computation 0.35 6% 0.07 6%
Budget Allocation 0.12 2% 0.02 2%
Negative Tuning Effects Mitigation
Concept Extraction 1.22 21% 0.26 22%
CCG Construction 0.71 12% 0.14 12%
Consistency Checking 0.34 6% 0.06 5%
Total Selection Time 5.83 100% 1.18 100%
Recovery Training (Full Data) 83.63 57.97
Recovery Training (PASER) 26.71 4.29
Time Saving 56.92 68.1% 53.68 92.6%
In the experiments, we follow the standard settings in the main body experiments: |B|=10.4K (20%) for Alpaca and |B|=103.2K (4%) for LaMini. Our component-wise time consumption analysis reveals the relative computational costs across PASER’s pipeline, as detailed in Table 18. The semantic-structural clustering process (S2RIC) consumes approximately 54% of the total selection time, with the Diffusion Kernel phase (30%) requiring the most computation in this component due to eigendecomposition operations.
The capability degradation assessment and budget allocation constitute about 8% of the selection time. Here, JSD computation is the most intensive operation (6%), while budget allocation using Equation 6 requires minimal overhead (2%). This aligns with our theoretical analysis, as JSD computation involves calculating probability distributions across the vocabulary space for each token.
The negative tuning effects mitigation component, particularly building and maintaining the Concept Consistency Graph, accounts for 39% of the selection process. Concept extraction (22%) and CCG construction during sample selection (12%) are the key operations. For Alpaca, the complete CCG-based filtering process required around 2.27 minutes but prevented negative tuning effects that would have otherwise caused performance degradation of 0.68-2.39 points across tasks.
For larger datasets like LaMini (2.58M samples), we implemented optimizations to maintain scalability: (1) approximate $k$-nearest neighbors with locality-sensitive hashing for manifold learning, reducing complexity from $O(N^{2})$ to $O(N\log N)$; (2) parallel processing for JSD computation across multiple CPU cores; and (3) incremental CCG updates with optimized data structures. With these optimizations, the complete data selection process for LaMini took approximately 1.18 hours, significantly less than the days required for full dataset recovery training.
Despite the overhead of data selection (5.83 minutes for Alpaca and 1.18 hours for LaMini), PASER achieves substantial overall time savings: 68.1% for Alpaca and 92.6% for LaMini compared to full dataset training. These savings stem from the reduced training data volume (20% for Alpaca and 4% for LaMini) and the focus on efficiency-driven sample selection that prioritizes samples with high information-to-computation ratios.
The empirical measurements confirm our theoretical analysis in Section 3.5 - with the number of concepts per sample $C$ being relatively small (average of 5-7 concepts per instruction), the dominant time complexity factor remains $O(N\log N)$. This enables PASER to scale efficiently to large instruction tuning datasets while maintaining its performance advantages over conventional recovery approaches.
Appendix M Case Study for Recovery Instruction Clustering
@@
M.1 Cluster 1: Basic Factual Knowledge and Information Retrieval
@@
M.2 Cluster 2: Language Understanding and Translation
@@
  • Instruction: “Difference in meaning between "done freely" and "freely done¨? For instance, is there any difference in meaning between these two sentences?”
@@
M.3 Cluster 3: Logical Reasoning and Problem Solving
@@
M.4 Cluster 4: Creative Writing and Text Generation
@@
M.5 Cluster 5: Summarization and Information Extraction
@@
  • Instruction: “Provide a brief summary of the article "A Brief History of the Automobile Industry".”
@@
Appendix N Visualization for Recovery Instruction Clustering
To better illustrate the semantic relationship of clusters after our Semantic-Structural Recovery Instruction Clustering, we provide the visualization results in the Figure 6. Notably, the numbered samples in each cluster correspond to those in above Appendix M. For example, the instructions numbered as 1,2,3 in Language Understanding cluster are the three samples provided in Appendix M.2. From this figure, we can find that our approach successfully identifies meaningful semantic structures in the instruction space that correspond to different LLM capabilities.
Figure 6: Visualization for the semantic-structural clustering of recovery instruction data by LLM capabilities. For ease of presentation, only the first two manifold dimensions are utilized.
Appendix O Case Study for Negative Tuning Effects Mitigation
To illustrate the negative tuning effects mitigation process in PASER, we present the following case study, including both initial concept consistency graph construction and new sample evaluation.
O.1 Constructing Initial Concept Consistency Graph
@@
O.2 Evaluating New Samples
We now evaluate two new samples using PASER’s negative tuning effects mitigation process.
O.2.1 Sample 1: Accepted
@@
Applying PASER’s negative tuning effects mitigation process:
@@
Applying PASER’s negative tuning effects mitigation process:
@@
O.3 Conclusion
This case study demonstrates PASER’s negative tuning effects mitigation process in action. By accepting Sample 1, PASER allows for the introduction of new, relevant concepts that expand the concept consistency graph without introducing conflicts. By rejecting Sample 2, PASER prevents the introduction of potentially misleading relationships between existing concepts from different domains, thus mitigating the risk of negative tuning effects during the recovery process.
Appendix P Implementation Details and Hyperparameter Settings
Most of experiments are conducted on the server with 8 GPUs (each with 48GB VRAM). For experiments involving the 70B model and MOE model, we use a high-performance GPU cluster where each GPU has 80GB VRAM. For the Semantic-Structural Recovery Instruction Clustering, we use consistent settings across all experiments: diffusion time $t$ is automatically selected using the spectral gap method, and the embedding dimension $d$ is set to 16. The optimal number of clusters $K$ is determined adaptively through NMF approximation error minimization, typically resulting in 8-12 clusters for Alpaca and 15-20 clusters for LaMini. For the JSD calculation in capability degradation score, we use a temperature $\tau$=1.0 for the output probability distribution. The computational cost is approximated using the quadratic term of sequence length with a coefficient of 1.0 across all experiments. For concept extraction, we use a maximum of 10 concepts per instruction-response pair with a minimum phrase length of 2 words and a maximum of 4 words. The concept similarity threshold for consistency checking is set to 0.75 across all experiments. We maintain these same hyperparameter settings across all models and pruning schemes to ensure fair comparison. During the recovery post-training phase, we take the the low-rank approximation, LoRA [Hu et al., ], to improve the efficiency. The corresponding hyperparameters are set as following: rank=8, batch size=64, epochs=2, learning rate = 1e-4 (Alpaca series experiments), 5e-5 (LaMini series experiments). As for the structured pruning, we set the pruning ratio as 25% for LLaMA2-7B/LLaMA3-8B/Baichuan2-7B and 50% for LLaMA2-13B/LLaMA2-70B/Baichuan-13B models. For the other two kinds of pruning schemes, we follow the previous work [Frantar and Alistarh, 2023]. Specifically, we adopt the 2:4 semi-structured sparsity patterns and implement 50% unstructured weight sparsity. Except the experiments for recovery post-training efficiency analysis, we set the ratio of recovery data budget $B$ to original dataset size $N$ as 20% for Alpaca and 4% for LaMini. As for the implementation of concept extraction in Section 3.4, we use the open-source library rake-nltk  [^6] https://pypi.org/project/rake-nltk/. To ensure statistical robustness, all the results reported in this paper are the averages of five runs with different seeds. Statistical significance is also assessed using two-tailed independent t-tests, with results considered significant when $p<0.01$. For facilitating the reproduction of our work, we provide the code in the supplementary materials, also seen in anonymous github https://anonymous.4open.science/r/PASER-E606.
Appendix Q Limitation Analysis
@@
Limitations of Concept Consistency Graph. In addition to the above analysis, we further discuss the concept consistency graph’s potential limitations. Indeed, while CCG helps mitigate negative tuning effect, we acknowledge there are scenarios where semantic conflicts might not be fully captured:
@@
Appendix R Ethics Statement